OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and perpendicular to y = -x + 6. y = -x + 6 y = x - 6 y = -x - 2 y = x + 2

OpenStudy (mathstudent55):

There is a relationship between the slopes of perpendicular lines. Do you know what it is?

OpenStudy (anonymous):

not at all im very bad at math but i have to do good on this

OpenStudy (mathstudent55):

The slopes of perpendicular lines are negative reciprocals. That means that when you multiply them you get -1.

OpenStudy (anonymous):

multiply there slopes...if answer is -1 then they are perp.

OpenStudy (anonymous):

how do i multiply slope

OpenStudy (mathstudent55):

The math term for slopes of perpendicular lines is that they are negative reciprocals. The way to find a slope of a perpendicular is to flip the fraction and change the sign. For example, if a line has slope 3/4, to find the slope of the perpendicular, you flip the fraction, getting 4/3, then you change the sign to -4/3.

OpenStudy (mathstudent55):

In your case, you are given the equation of a line as y = -x + 6 The slope is the coefficient (the number multiplying) the x variable. -x means -1 * x, so the slope of the given line is -1. Ok so far?

OpenStudy (mathstudent55):

Now you need to slope of the perpendicular. Flip -1. -1 is the same as -1/1. Flip it and you still have -1/1. Now change the sign to 1/1 which is the same as 1. That means the slope of the perpendicular is 1. That means we will have an equation of the form y = x + b

OpenStudy (mathstudent55):

To find b, we use the given point in for x and y, and we solve for b. y = x + b -2 = -4 + b Solve for b: 2 = b, or b = 2 That means the equation of the perpendicular is y = x + 2

OpenStudy (anonymous):

is it b?